OpenStudy (madeleine2):

Write the slope-intercept form of the line that has a slope of 2 and intersects the line, 2x - 3y = 6 at x = 3

OpenStudy (pythagoras123):

General formula of a line: y = mx+c Where m is the gradient and c is the y-intercept. A line with a slope of 2 has the formula: y = 2x + c -------------------------(1) We can re-arrange the equation 2x-3y=6 to get: 3y = 2x - 6 Sub x = 3 into this equation to get: y = 0 Hence the unknown line passes through the point (3,0). Sub this point into equation (1) to get: 0 = 6 + c c = -6 Hence the equation of the line is y = 2x - 6.